OpenStudy (anonymous):

find a function f(x) and a number a such that 4 + integral from a to x^(2) f(t)/t dt = 2sqt x

myininaya (myininaya):

\[4+\int\limits_{a}^{x^2}\frac{f(t)}{t} dt =2 \sqrt{x} ?\]

myininaya (myininaya):

let's try something \[\text{ Let } g(t) =\frac{f(t)}{t} \text{ let G(t) be antiderivative} \] \[4 +\int\limits_{a}^{x^2}g(t) dt =2 \sqrt{x}\] \[4+G(t)|_a^{x^2} =2 \sqrt{x}\] \[4+G(x^2)+G(a)=2 \sqrt{x}\] now if we differentiate both sides we get \[0+(x^2)' g(x^2)+0=\frac{1}{\sqrt{x}}\] \[2xg(x^2)=\frac{1}{\sqrt{x}} => g(x^2)=\frac{1}{ \sqrt{x}} \cdot \frac{1}{2x}\] \[g(x^2)=\frac{1}{2x \sqrt{x}}\] \[\frac{f(x^2)}{x^2}=\frac{1}{2 x \sqrt{x}}\] \[f(x^2)=\frac{x^2}{2 x \sqrt{x}}\] \[f(x^2)=\frac{x^2}{2 x^{1+\frac{1}{2}}}=\frac{x^2}{2 x^{\frac{3}{2}}}\] \[f(x^2)=\frac{(x)^2}{2 ((x)^2)^\frac{1}{2} \cdot ((x)^2)^\frac{1}{4}}\]

myininaya (myininaya):

\[f(x)=\frac{x}{2 x ^\frac{1}{2} \cdot x ^\frac{1}{4}}\]

myininaya (myininaya):

so \[f(x)=\frac{x}{2 x^\frac{3}{4}}=\frac{1}{2}x^{1-\frac{3}{4}}=\frac{1}{2}x^{\frac{1}{4}}\]

myininaya (myininaya):

\[4+\int\limits_{a}^{x^2} \frac{1}{2} \cdot \frac{t^\frac{1}{4}}{t} dt =2 \sqrt{x}\] \[4+\frac{1}{2} \int\limits_{a}^{x^2}t^{\frac{-3}{4}} dt = 2 \sqrt{x}\] \[4+\frac{1}{2} \cdot \frac{t^{\frac{-3}{4}+1}}{\frac{-3}{4}+1}|_a^{x^2}=2 \sqrt{x}\] \[4+\frac{1}{2} \cdot \frac{t^\frac{1}{4}}{\frac{1}{4}}|_a^{x^2}=2 \sqrt{x}\] \[4+\frac{1}{2} \cdot 4[(x^2)^\frac{1}{4}-a^\frac{1}{4}]=2 \sqrt{x}\] \[4+2[x^\frac{1}{2}-a^\frac{1}{4}] =2 \sqrt{x}\] \[2 \sqrt{x} +4 -a^\frac{1}{4}= 2 \sqrt{x}\]

myininaya (myininaya):

\[=> 4- a^\frac{1}{4}=0\]

myininaya (myininaya):

solve for a

myininaya (myininaya):

\[4 =a^\frac{1}{4}\]

myininaya (myininaya):

\[4^4=a\]

myininaya (myininaya):

this is a cute problem

myininaya (myininaya):

i should give this to my cal students

OpenStudy (zarkon):

you have a small typo...

myininaya (myininaya):

no!

OpenStudy (zarkon):

after \[4+2[x^\frac{1}{2}-a^\frac{1}{4}] =2 \sqrt{x}\]

myininaya (myininaya):

darn it stupid 2

myininaya (myininaya):

\[4-2 a^\frac{1}{4}=0\]

myininaya (myininaya):

\[4=2 a^\frac{1}{4}\]

myininaya (myininaya):

\[2=a^\frac{1}{4}\]

myininaya (myininaya):

\[2^4=a\]

OpenStudy (zarkon):

nice problem

myininaya (myininaya):

did i go the long way?

OpenStudy (zarkon):

not that I can see

myininaya (myininaya):

ok it does seem to be lengthy it should be bonus they probably won't figure out anyways

OpenStudy (zarkon):

prob not...probably get stuck at the f(x^2) part ... nice work

myininaya (myininaya):

yes thats what i was thinking